Polar form of i to the power 25 whole to the power 3

Grade:11

1 Answers

Vikas TU
6 years ago
We can write,
z = (i^25)^3 => i^75 => =i^3n = -i
=> -i = 0 – i
 
=> Polar Form of z  = r ( cos a  + i sin a).
=> cos(-pi/2) + isin(-pi/2)
=> cos(pi/2) - isin(pi/2)